OpenStudy (anonymous):

What is the solution of the equation? \[-10+\sqrt{x +8=-4}\]

OpenStudy (solomonzelman):

\[-10+\sqrt{x+8}=-4~~~~~~~like~~this~?\]

OpenStudy (solomonzelman):

\[-10+\sqrt{x+8}=-4\]\[\sqrt{x+8}=6\]\[x+8=36\] take it from here, and ask about any step that you don't get...

OpenStudy (anonymous):

okay I see what you did with the first step but how did you get rid of the square root and get 36?

OpenStudy (solomonzelman):

I raised each side to the second power.

OpenStudy (anonymous):

okay I get it now. would I add or subtract 8 from each side now?

OpenStudy (anonymous):

x=28?